OpenStudy (anonymous):

I NEED THIS DONE ASAP!!!!! HELP ME PLZ!!!!! Which property is shown? -2 + (-15) = -(2 + 15) identity property associative property commutative property opposite of a sum property What is the solution to the inequality? 5x + 8 < 12 x > -4 x < -4 x > 4 x < 4 What is the solution to the equation? 2(x + 7) = 10x - 10 x = 3 x = 2 x = -2 x = -3 What is 100,000,000 written as a power of ten?

OpenStudy (pixiedust1):

Okay, I know for "What is the solution to the equation" What we wanna do is make both sides of the equation balanced. Therefore, we take one of the answer options and substitute it in for x

OpenStudy (pixiedust1):

Let's try x=3 2(3+7) = 10*3-10 3+7=10 10*3=30 2*10=20 30-10=20 Success! With x=3, we get that 3=3 Therefore, x=3 is the answer

OpenStudy (anonymous):

Thank you thank you so much!!

OpenStudy (pixiedust1):

For the last one, a power of 10 means like 10 with an exponent. So, 10^n will equal 100,000,000 Just to make this quick, let's try 8 10^8=100,000,000

OpenStudy (anonymous):

OK thank you thank you

OpenStudy (pixiedust1):

So, its 1.0*10^8

OpenStudy (the_fizicx99):

Um no not really, if you didn't have the options you wouldn't be able to solve. But you can still solve, that method is considered cheating to a majority of teachers @PixieDust1. 2(x + 7) = 10x - 10 2x +14 =10x -10 -14 -14 2x = 10x -24 -10x -10x -8x = -24 /-8 /-8 x = 3

OpenStudy (pixiedust1):

Yeah, that's the long way. But they both give the same answer

OpenStudy (pixiedust1):

It's just a shortcut. I'm sure her lesson taught her the long way.

OpenStudy (pixiedust1):

And no, no method is considered cheating.

OpenStudy (the_fizicx99):

That's still considered cheat fyi. You're not suppose to use the options.. but who am I to correct you :/

OpenStudy (anonymous):

Oh I am sorry that is my bad

OpenStudy (anonymous):

Aplogizes on my part

OpenStudy (pixiedust1):

It's okay lila, there's more then one way to solve things. :)

OpenStudy (anonymous):

:)

OpenStudy (the_fizicx99):

You're not suppose to use the option.. its a shortcut, yes. Normally solving like that is how you would arrive at 3. My point is if you didn't have the option, how would you solve?

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

Good that all of this is settled! :)

OpenStudy (pixiedust1):

Yup! Now we can *hopefully* return to the question! My apologies! Okay, so I guess the only two left are the first two questions

OpenStudy (anonymous):

Yup!:)

OpenStudy (anonymous):

Yes defintinly!

OpenStudy (anonymous):

I need the first two questions anwsered please.

OpenStudy (anonymous):

the answer to the first one is commutative

OpenStudy (anonymous):

Thanks

OpenStudy (pixiedust1):

@shelbygt520 Can you explain why? I wasn't sure on this one

OpenStudy (anonymous):

Anytime. the answer to the second one is 5x + 8 > -12 5x + 8 - 8 > -12 - 8 5x > -20 5x/5 > -20/5 x > -4 Hope this helps! http://answers.yahoo.com/question/index?qid=20110309165523AAn0uje

OpenStudy (anonymous):

Thanks!!

OpenStudy (anonymous):

No problem do you need the third one?

OpenStudy (anonymous):

Nope I am good thank you

OpenStudy (anonymous):

Anytime :)